OpenStudy (anonymous):

what are the factors of 25t^2-1

OpenStudy (anonymous):

you could use the quadratic equation \[-b \pm \sqrt{b ^{2}-4ac } \over 2a\]

OpenStudy (anonymous):

y = ax + b + c

OpenStudy (anonymous):

\[ a^2 - b^2 =(a-b)(a+b)\\ 25 t^2 -1 =(5t-1)(5t+1) \]

OpenStudy (anonymous):

^Yeah, use the difference of two squares method when you can.